Suppose you are using total internal reflection to make an efficient corner reflector. If there is air outside and the incident angle is 54°, what must be the minimum index of refraction of the material from which the reflector is made?

answer to 4 decimal places.


1
Expert's answer
2021-06-01T14:35:57-0400

Answer:-


"Sin(\theta_c)=\frac{n_2}{n_1}\\
n_1=\frac{1.00293}{sin54^o}\\
\boxed{n_1=1.2364}"

the minimum refractive index = 1.2364